# MADGEN - Mass-Spec attends to De Novo Molecular generation

### v1

@@ Abstract
The annotation (assigning structural chemical identities) of MS/MS spectra remains a significant challenge due to the enormous molecular diversity in biological samples and the limited scope of reference databases. Currently, the vast majority of spectral measurements remain in the “dark chemical space” without structural annotations. To improve annotation, we propose MADGEN (Mass-spec Attends to De Novo Molecular GENeration), a scaffold-based method for de novo molecular structure generation guided by mass spectrometry data. MADGEN operates in two stages: scaffold retrieval and spectra-conditioned molecular generation starting with the scaffold. In the first stage, given an MS/MS spectrum, we formulate scaffold retrieval as a ranking problem and employ contrastive learning to align mass spectra with candidate molecular scaffolds. In the second stage, starting from the retrieved scaffold, we employ the MS/MS spectrum to guide an attention-based generative model to generate the final molecule. Our approach constrains the molecular generation search space, reducing its complexity and improving generation accuracy. We evaluate MADGEN on three datasets (NIST23, CANOPUS, and MassSpecGym) and evaluate MADGEN’s performance with a predictive scaffold retriever and with an oracle retriever. We demonstrate the effectiveness of using attention to integrate spectral information throughout the generation process to achieve strong results with the oracle retriever.
@@ 3.1 Scaffold Retrieval
We formulate the scaffold retrieval as a ranking problem. Given a set of scaffold candidates $\mathbb{S}$, the goal is to use a neural network to score each candidate $S\in\mathbb{S}$ given $X$ such that scaffold with highest score $S^{*}$ maximally resembles the correct scaffold $S^{\mathrm{gt}}$.
@@
After training, we access the score of each scaffold candidate via cosine similarity between the mass spectra embedding $f_{X}(X)$ and the scaffold embedding $f_{S}(S)$. The scaffolds with the top-k cosine similarity scores are selected and used as the starting points for stage 2 of the molecular generation process.
@@ 4.2 Experimental Setup and Evaluation Metrics
The model was trained using a graph transformer with 5 layers and 50 diffusion steps. We employed the AdamW optimizer with a learning rate of $1\times 10^{-5}$. Full training details and hyperparameters can be found in Appendix A.2.
For candidate pool selection, the following approaches were employed:
@@
  • MassSpecGym: the candidate pool consists of 256 molecules for each test molecule. These candidates are selected based on the molecular formula provided by the MassSpecGym dataset.
@@
  • Top-k accuracy: Measures the likelihood of the generating the true target structure among the top-k generated molecules. We report the results for k=1,10
@@ 4.3 Results
Our experiments, summarized in Table 2 , evaluate model performance on three datasets: NIST23, CANOPUS, and MassSpecGym, using both predictive and oracle retrievers. For the scaffold prediction task, we report a Scaffold Prediction Accuracy (SPA) for the predictive retriever ranging from 34.8% to 57.8%. Notably, the NIST23 dataset achieves the highest SPA of 57.8%, reflecting its lower scaffold diversity compared to CANOPUS and MassSpecGym, which have more complex scaffolds.
The metrics for the scaffold-based generation task reveal that the low scaffold prediction accuracy of the predictive retriever constrains molecular generation performance. For instance, on the NIST23 dataset, the predictive retriever yields a top-1 accuracy of 10.3%, while for CANOPUS and MassSpecGym, the top-1 accuracies are 1.0% and 0.8%, respectively. Despite these challenges, the predictive retriever demonstrates moderate performance improvements compared to baseline methods like Spec2Mol and random generation.
In contrast, the oracle retriever, which has access to the correct scaffold, dramatically boosts performance. On NIST23, MADGEN achieves a top-1 accuracy of 49% and a top-10 accuracy of 65.5%, demonstrating the model’s capacity to generate accurate molecular structures if the scaffold is known. Similarly, when using the oracle retriever, the performance on CANOPUS and MassSpecGym is significantly higher than the predictive retriever, with top-1 accuracies of 8.9% and 18.8%, respectively, showing the clear advantage of having access to correct scaffold information. Importantly, MADGEN outperforms the best published state-of-the-art performance (last row in Table 2) reported for the MassSpecGym dataset (Bushuiev et al., 2024) when using random chemical generation. The high top-1 and Top10 accuracy for the NIST23 dataset can be attributed to its smaller number of free atoms. MADGEN’s task of completing the target molecule by adding edges to these free atoms is easier with a smaller number of free atoms. CANOPUS has the highest number of average free atoms, and the lowest top-1 and top-10 accuracies.
@@
Table 2: Performance metrics for various datasets using both predictive and oracle retrievers. The table presents top-1 and top-10 accuracy, Tanimoto similarity, and Maximum Common Edge Substructure (MCES) scores. Best performance for each dataset is bold. The second best performance for each dataset is underline
@@ 5 Conclusion & Future Work
De novo annotation of mass spectrometry data is notoriously difficult, with a current best accuracy of 0% on the MassSpecGym dataset. MADGEN offers a novel two-stage framework for spectra-guided de novo annotation. The first stage, scaffold retrieval, is a new problem formulation whose solution can provide partial insight in regard to the molecular backbone of the measured spectra. Such insights may shed light on the molecule’s class or properties. Our results show that this problem is challenging, achieving a scaffold prediction accuracy of 34.8%-57.8% for the three datasets. The second stage, de novo generation from an existing scaffold showed excellent results when using an oracle scaffold predictor, achieving an accuracy of 8.9%-49% across the three dataset. For the MassSpecGym benchmark, we were able to achieve an accuracy of 0.8%, an almost 50% increase in accuracy over MassSpecGym baseline. As with other tools, e.g., (Goldman et al., 2023), we conjecture that performance of MADGEN can be increased by incorporate additional data in the form of peak chemical formulae or molecular properties that correlate with fragmentation patterns. Potentially, the scaffold problem can be made easier if larger more distinct scaffold structures were utilized instead of the Murcko scaffold used herein. A bigger scaffold can in turn facilitate the de novo generation task. Further, an end-to-end MADGEN may reduce the compounding of errors across the two stages.
@@ A.2 Training Hyperparameters
The model is trained with a batch size of 64 and employed 47 workers for data loading. The learning rate is set to $1\times 10^{-5}$, while weight decay is configured at $1\times 10^{-12}$. Training proceeds for 1000 epochs, with the model logging progress every 40 steps.
A Markov bridge process with 50 steps is employed during training, and a cosine noise schedule is employed.

### v3

@@ Abstract
The annotation (assigning structural chemical identities) of MS/MS spectra remains a significant challenge due to the enormous molecular diversity in biological samples and the limited scope of reference databases. Currently, the vast majority of spectral measurements remain in the “dark chemical space” without structural annotations. To improve annotation, we propose MADGEN (Mass-spec Attends to De Novo Molecular GENeration), a scaffold-based method for de novo molecular structure generation guided by mass spectrometry data. MADGEN operates in two stages: scaffold retrieval and spectra-conditioned molecular generation starting with the scaffold. In the first stage, given an MS/MS spectrum, we formulate scaffold retrieval as a ranking problem and employ contrastive learning to align mass spectra with candidate molecular scaffolds. In the second stage, starting from the retrieved scaffold, we employ the MS/MS spectrum to guide an attention-based generative model to generate the final molecule. Our approach constrains the molecular generation search space, reducing its complexity and improving generation accuracy. We evaluate MADGEN on three datasets (NIST23, CANOPUS, and MassSpecGym) and evaluate MADGEN’s performance with a predictive scaffold retriever and with an oracle retriever. We demonstrate the effectiveness of using attention to integrate spectral information throughout the generation process to achieve strong results with the oracle retriever. Our code is available at https://github.com/HassounLab/MADGEN
@@ 3.1 Scaffold Retrieval
We formulate the scaffold retrieval as a ranking problem. Given a set of scaffold candidates $\mathbb{S}$, the goal is to use a neural network to score each candidate $S\in\mathbb{S}$ given $X$ such that scaffold with highest score $S^{*}$ maximally resembles the correct scaffold $S^{\mathrm{gt}}$. We rank on a candidate set from which the target molecules have been removed, introducing the possibility that the true scaffold may not be present in the set.
@@
After training, we access the score of each scaffold candidate via cosine similarity between the mass spectra embedding $f_{X}(X)$ and the scaffold embedding $f_{S}(S)$. To improve scaffold ranking accuracy, we introduce a frequency-based aggregation approach. For each data point, we first retrieve the top-$k$ ranked candidate scaffolds. The frequency of each scaffold appearing in these top candidates is then computed, and the most frequently occurring scaffold for each formula is selected as the predicted scaffold. This method refines scaffold selection by leveraging consensus among top-ranked candidates, leading to improved scaffold prediction accuracy (SPA).
@@ 4.2 Experimental Setup and Evaluation Metrics
The model was trained using a graph transformer with 5 layers and 50 diffusion steps. We employed the AdamW optimizer with a learning rate of $1\times 10^{-5}$. Full training details and hyperparameters can be found in Appendix A.2. For candidate pool selection, the following approaches were employed:
@@
  • MassSpecGym: the candidate pool consists of 256 molecules for each test molecule. These candidates are selected based on the molecular formula provided by the MassSpecGym dataset. The target molecule is removed from the candidate pool.
@@
  • Top-k accuracy: Measures the likelihood of the generating the true target structure among the top-k generated molecules. We report the results for k=1,10. The generated molecules are ranked by the probabilistic nature of the model.
@@ 4.3 Results
Our experiments, summarized in Table 2 , evaluate model performance on three datasets: NIST23, CANOPUS, and MassSpecGym, using both predictive and oracle retrievers. For the scaffold prediction task, we report a Scaffold Prediction Accuracy (SPA) for the predictive retriever ranging from 13.2% to 40.3%. Notably, the NIST23 dataset achieves the highest SPA of 40.3%, reflecting its lower scaffold diversity compared to CANOPUS and MassSpecGym, which have more complex scaffolds.
The metrics for the scaffold-based generation task reveal that the low scaffold prediction accuracy of the predictive retriever constrains molecular generation performance. For instance, on the NIST23 dataset, the predictive retriever yields a top-1 accuracy of 4.6%, while for CANOPUS and MassSpecGym, the top-1 accuracies is 2.10 % and 1.31%. Despite these challenges, the predictive retriever demonstrates moderate performance improvements compared to baseline methods like Spec2Mol and random generation.
In contrast, the oracle retriever, which has access to the correct scaffold, dramatically boosts performance. On NIST23, MADGEN achieves a top-1 accuracy of 49.0% and a top-10 accuracy of 65.5%, demonstrating the model’s capacity to generate accurate molecular structures if the scaffold is known. Similarly, when using the oracle retriever, the performance on CANOPUS and MassSpecGym is significantly higher than the predictive retriever, with top-1 accuracies of 18.7% and 10.5%, respectively, showing the clear advantage of having access to correct scaffold information. Importantly, MADGEN outperforms the best published state-of-the-art performance (last row in Table 2) reported for the MassSpecGym dataset (Bushuiev et al., 2024) when using random chemical generation. The high top-1 and Top10 accuracy for the NIST23 dataset can be attributed to its smaller number of free atoms. MADGEN’s task of completing the target molecule by adding edges to these free atoms is easier with a smaller number of free atoms. CANOPUS has the highest number of average free atoms, and the lowest top-1 and top-10 accuracies.
@@
Table 2: Performance metrics for various datasets using both predictive and oracle retrievers. The table presents top-1 and top-10 accuracy, Tanimoto similarity, and Maximum Common Edge Substructure (MCES) scores. Best performance for each dataset is bold. The second-best performance for each dataset is underlined
@@ 5 Conclusion & Future Work
De novo annotation of mass spectrometry data is notoriously difficult, with a current best accuracy of 0% on the MassSpecGym dataset. MADGEN offers a novel two-stage framework for spectra-guided de novo annotation. The first stage, scaffold retrieval, is a new problem formulation whose solution can provide partial insight in regard to the molecular backbone of the measured spectra. Such insights may shed light on the molecule’s class or properties. Our results show that this problem is challenging, achieving a scaffold prediction accuracy of 13.2%-40.3% for the three datasets. The second stage, de novo generation from an existing scaffold showed excellent results when using an oracle scaffold predictor, achieving an accuracy of 10.5%-49% across the three dataset. For the MassSpecGym benchmark, we achieved an accuracy of 2.10% and 1.31%. As with other tools, e.g., (Goldman et al., 2023), we conjecture that performance of MADGEN can be increased by incorporating additional data in the form of peak chemical formulae or molecular properties that correlate with fragmentation patterns. Potentially, the scaffold problem can be made easier if larger more distinct scaffold structures were utilized instead of the Murcko scaffold used herein. A bigger scaffold can in turn facilitate the de novo generation task. Further, an end-to-end MADGEN may reduce the compounding of errors across the two stages.
Acknowledgments
Research reported in this publication was supported by the National Institute of General Medical Sciences of the National Institutes of Health under award number R35GM148219. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
@@ A.2 Training Hyperparameters
The model is trained with a batch size of 64 and employed 47 workers for data loading. The learning rate is set to $2\times 10^{-4}$, while weight decay is configured at $1\times 10^{-12}$. Training proceeds for 2000 epochs, with the model logging progress every 40 steps.
A Markov bridge process with 100 steps is employed during training, and a cosine noise schedule is employed.